how to determine the vertex and zeros of the equation y= 1/4(x+2)^2-5
Answer by stanbon(75887)

how to determine the vertex and zeros of the equation y= 1/4(x+2)^2-5
----
The vertex in that form is (-2,-5)
-----
Zeros ??
If x = 0, y = 1-5 = -4
----
Solve:
(1/4)(x+2)^2-5 = 0
(x+2)^2 = 20
x+2 = +-2sqrt(5)
x = -2+2sqrt(5) or x = -2-2sqrt(5)
---------------------------
Cheers,
Stan H.
---------------------